# It is Simple Sometimes: A Study On Improving Aspect-Based Sentiment Analysis Performance

### v1

@@ 1 Introduction
We make our code publicly available to ensure reproducibility and foster future research. [^1] https://github.com/lautel-sony/PFInstruct
@@ 4.3 Domain generalization
As expected, we observe a drop in performance compared to training in-domain, which is especially large in ATE. When training on Lapt14 and evaluating each restaurant dataset, the strategy of adding a noisy prefix achieve the highest F1-scores on average across tasks, except for ATE Rest16. When training in restaurant domain and evaluating in Lapt14 we observe a similar trend in performance decrease, except for AOOE. Interestingly in this task, PFInstruct-Noise suffer a performance drop of $-3.35$ F1 (compared to PFInstruct-Noise in Table 3) but the other model variants surpass their respective in-domain results with a major improvement of $+3.74$ F1 in PFInstruct-NER.
While the strategy of adding a noisy prefix seem beneficial to out-of-domain data performance too, with the exceptions mentioned above, looking closer we observe that the drop in performance when evaluated on out-of-domain data (compared to in-domain) is larger for PFInstruct-Noise models. Therefore, these results suggest that an NLP task prefix makes the model more robust to domain shifts.
@@
Figure 2: Out-of-domain evaluation. Models are trained (‘tr’) on one domain and evaluated (‘tst’) on a distinct domain. Legends indicate the prefix prompt used. ‘No’ stands for no use of prefix. RE is not evaluated for ATE (see Section 4).

### v2

@@ 1 Introduction
We make our code publicly available to ensure reproducibility and foster future research. [^1] https://github.com/lautel/PFInstruct
@@ 4.3 Domain generalization
As expected, we observe a general drop in performance compared to training in-domain, which is especially large in ATE. However, when training in restaurant domain and evaluating in Lapt14 for AOOE –see Figure 2 (d)–, all model variants surpass their respective in-domain results with a major improvement of $+3.53$ F1 in PFInstruct-NER. We conclude that the addition of more training data is beneficial for this task.
The addition of any kind of prefix helps to make the model more robust to out-of-domain data for AOOE, while it does not significantly hurt performance for ATSC. Interestingly in this task, the large variance shown by the model without prefix is reduced by the addition of a prefix, especially with RE and NER prefixes –see Figure 2 (c), (d)–.
While the strategy of adding a noisy prefix seem beneficial to out-of-domain data performance too, looking closer we make two observations: i) PFInstruct-Noise models show large variance regardless of the domain, and ii) the drop in performance when evaluated on out-of-domain data (compared to in-domain) is larger for PFInstruct-Noise models. Therefore, these results suggest that an NLP task prefix makes the model more robust to domain shifts.
@@
Figure 2: Out-of-domain evaluation. F1-scores are averaged across five random initialization seeds; error bars show the standard deviation. Models are trained (‘tr’) on one domain and evaluated (‘tst’) on a distinct domain. Legends indicate the prefix prompt used. ‘No’ stands for no use of prefix. RE is not evaluated for ATE (see Section 4).